OpenStudy (anonymous):

What is the volume of a cylinder with the following dimensions: Radius 6ft Height 24ft

OpenStudy (opcode):

\(\text{Cylinder formula:}\) \(volume = π \times r^2 \times h\)

OpenStudy (opcode):

Can you plug in the values?

OpenStudy (opcode):

Okay I'll do it for you one second ^_^.

OpenStudy (opcode):

\(volume = π \times 6^2 \times 24\)

OpenStudy (anonymous):

lol thank you. I'm in the middle of an exam. >.<

OpenStudy (opcode):

Okay from that we get: 2712.96